Mathematics 133 – Intensive Calculus for Science 1 Quiz 7 – November 11, 2005

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. No Calculators allowed on this quiz.

A) Compute f′(x) for each of the following (you do not need to simplify)

1) f(x) = ln(x^{2} + 3 sin(x))

2) f(x) = e^{x} arctan(x)

B) Find all critical points of f(x) = 2x^{3} − 3x^{2} + 1 and classify each as a local maximum, local minimum, or neither.